# Reliable one-bit quantization of bandlimited graph data via single-shot noise shaping

### v1

@@ 5 Discussion
Limitations. In our study, we focused on the brick-wall filter as a prototype of low-pass filters. Various other low-pass filters exist, such as (Defferrard et al., 2016; Chang et al., 2021; Nt and Maehara, 2019). We furthermore provided no robustness analysis of our method on approximately $r$-bandlimited data, and only use the graph incoherence to characterize the influence of the graph topology. As our experiment in Figure 3 (f)–(h) suggests, a more refined analysis might be needed to fully explain the observed quantization error on complex graphs for strongly bandlimited signals, cf. Experiment 1 in Section 4.
Impact Statement
This paper aims to advance the field of machine learning, in particular by improving the theoretical understanding of one-bit quantization of graph-structured data as a tool for representing and processing structured graph signals. While there are many potential societal consequences of our work, we do not foresee specific impacts that require separate discussion here.
Supplementary Material
We provide pseudo-code of the accelerated preprocessing algorithm of Meyer (2024) in Algorithm 3. It improves over Algorithm 1 by extracting kernel vectors $\bm{b}$ in a more efficient way. To this end, it splits the columns of $\bm{X}$ into blocks of size $r$ and iteratively extracts collections of $r$ kernel vectors per block. By slightly modifying the single kernel vectors in $r$ inner iterations, they can all be used before a new collection has to be generated. Generating $r$ kernel vectors is of the same complexity as generating a single kernel vector. Asymptotically, each of the $\lceil N/r\rceil$ outer iterations of Algorithm 3 thus requires as many operations as each of the $N$ iterations of Algorithm 1, leading to an overall complexity improvement of a factor $r$.
@@
Algorithm 3 Vector preprocessing (Meyer, 2024)

### v2

@@ 5 Discussion
Limitations. In our study, we focused on the brick-wall filter as a prototype of low-pass filters. Various other low-pass filters exist, such as Defferrard et al. (2016); Chang et al. (2021); Nt and Maehara (2019). Given that the brick-wall filter is the standard example of an idealized low-pass filter, and that to the best of our knowledge this is the first approach to graph quantization that allows one-bit quantization and comes with rigorous error guarantees in the $\mathcal{Q}\mathcal{E}_{L}$-measure, we believe that such a restriction is more than legitimate. We furthermore provided no robustness analysis of our method on approximately $r$-bandlimited data, and only use the graph incoherence to characterize the influence of the graph topology. As our experiment in Figure 3 (f)–(h) suggests, a more refined analysis might be needed to fully explain the observed quantization error on complex graphs for strongly bandlimited signals, cf. Experiment 1 in Section 4.
Appendix A Supplementary material
A.1 Efficient implementation of Algorithm 1
We provide pseudo-code of the accelerated preprocessing algorithm of Meyer [2024] in Algorithm 3. It improves over Algorithm 1 by extracting kernel vectors $\bm{b}$ in a more efficient way. To this end, it splits the columns of $\bm{X}$ into blocks of size $r$ and iteratively extracts collections of $r$ kernel vectors per block. By slightly modifying the single kernel vectors in $r$ inner iterations, they can all be used before a new collection has to be generated. Generating $r$ kernel vectors is of the same complexity as generating a single kernel vector. Asymptotically, each of the $\lceil N/r\rceil$ outer iterations of Algorithm 3 thus requires as many operations as each of the $N$ iterations of Algorithm 1, leading to an overall complexity improvement of a factor $r$.
@@ A.1 Efficient implementation of Algorithm 1
Algorithm 3 Vector preprocessing Meyer [2024]
@@
A.2 Additional numerical experiments
Experiment 3: Empirical and Theoretical Bit-Depth Scaling. In this experiment, we investigate the effect of quantization bit-depth on the reconstruction accuracy of bandlimited graph signals under the SSNS method, and compare its empirical performance with our theoretical error bounds. We consider the same graph topologies as in Experiment 1. For each graph, we generate bandlimited graph signals supported on the first $r=200$ graph Fourier modes, and generate $50$ independent signal realizations. We measure the relative squared reconstruction error (4.1) and report the average error over all realizations for each bit-depth.
Experiment 4: Wall-clock Runtime Comparison. We evaluate the wall-clock runtime of the two quantization strategies on representative graph datasets, namely grid, bunny, and swiss_roll. For each graph, we generate bandlimited signals with a fixed bandwidth of $r=50$ and perform quantization using the two approaches SSNS and SSS-R. The experiments are presented in Figure 6. All experiments are repeated over $5$ independent runs, and we report the average runtime.
We explicitly separate the computational cost of the graph Fourier basis (eigendecomposition) from the quantization step. Figure 6 shows that the quantization costs of SSS-R (excluding the eigendecomposition) are notably lower than the ones of SSNS. However, for large graphs, the eigenbasis computation dominates the total runtime of both algorithms.
Figure 6: Wall-clock runtime comparison for different graph types (1-bit quantization). Each bar shows the total runtime, decomposed into eigenbasis computation (blue) and quantization (orange).
Experiment 5: Graph-Signal Halftoning on 3D Meshes. Finally, we evaluate the proposed SSNS algorithm for halftoning of graph signals on 3D shapes and compare it with Sigma-Delta-Weight (SDW) in Krahmer et al. [2023] and memoryless scalar quantization (MSQ); see also related work on 3D halftoning in Gooran and Abedini [2020, 2022], Mao et al. [2017], Lou and Stucki [1998], Abedini et al. [2023], Abedini [2023].
This experimental design follows the classical motivation of digital halftoning: although the displayed signal is binary, perceptual integration by the human visual system acts as an implicit low-pass filter, so visual appearance is primarily determined by low-frequency content Floyd and Steinberg [1976], Lyu and Wang [2023], Krahmer and Veselovska [2022]. While the signal is not bandlimited and therefore lies outside the scope of our theoretical analysis, the three methods produce visually distinct halftoning patterns, illustrating different ways of shaping quantization artifacts on graphs. We emphasize that with higher bandwidth $r$, among the compared approaches, only SSNS is supported by provable error bounds, highlighting its role as a theoretically justified alternative for graph-based halftoning.
(a) Original bunny shape
(c) Bunny shape, 1 bit SDW, $r=20$
(d) Bunny shape, 1 bit SSNS, $r=20$
(b) Bunny shape, 1 bit MSQ
(e) Bunny shape, 1 bit SDW, $r=50$
(f) Bunny shape, 1 bit SSNS $r=50$
Figure 7: Halftoning of the Stanford Bunny mesh graph signal using 1-bit quantization. (a) Original signal. (b) Memoryless scalar quantization (MSQ). (c),(e) Sigma-Delta-Weight (SDW) Krahmer et al. [2023] with different bandwidth $r$. (d),(f) Proposed Single-Shot Noise Shaping (SSNS) with different different bandwidth $r$. All halftoned outputs are binary signals $\bm{q}\in\{-1,1\}^{N}$ visualized on the mesh.